• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Glashausser

    Log entry time 16:05:46 on January25,2002

    Entry number 80773

    keyword=Day shift log

    8AM In controlled access for calorimeter angle change for kinematics 2B.
    Calorimeter set to 56deg at 7.1 m.
    HRS L moved to 28.281 deg.

    9:30 Start running on optics tgt (Total time elapsed since call for access to
    change calorimeter position was 2 hrs.)

    10:05 Beam down
    10:39 beam back
    10:49 start run 1643 production 2Bb
    11:25 Sweep magnet on for 2Bd
    12:05 Finish 2Bd
    Move spectro to 29.824 deg
    Change tgt to LH2 Rad
    Have to lower current to about 8 to get T1 down below 1M and T5 to 1.3 kHz
    12:40 have to reboot cryotgt
    Discuss whether we really have to stop runs after 290K. Run length now
    5-6 min at 8microA. Can take 12 mics probably. Ting says too many
    problems in espace with split runs.
    12:55 cryo reboot completed (took unusually long time) Had to reboot CODA
    to start next run (1650)
    13:25 Ting says he got 400 RCS events in one 290K run in this kinematics (1647).
    We have done 5 of these. He says to do one more to be safe.
    13:31 Turn sweep magnet power down and off
    13:32 Start 2Bf
    14:08 Reboot CODA
    14:20 Finish 2Bf Change to optics target
    14:28 Start 2Bg (Optics)
    14:42 Finish 2B kinematics
    Tech here to start survey
    15:21 Back from Hall after changing calorimeter angle to 79 deg, 5.2 m
    2C trigger downloaded. Threshold changed to 300MeV
    15:27 Power Permit
    Spectro L being reset. Angle 19.134 deg. Waited to move until
    arm out of Hall.)
    15:41 Beam in Hall (Less than one hr!)
    15:42 Start KIN 2C

    Bottom Line: KIN 2B completed. Apart from a few target reboots, no
    particular problems.